# UniMODE: Unified Monocular 3D Object Detection

### v4

@@ Abstract
Realizing unified monocular 3D object detection, including both indoor and outdoor scenes, holds great importance in applications like robot navigation. However, involving various scenarios of data to train models poses challenges due to their significantly different characteristics, e.g., diverse geometry properties and heterogeneous domain distributions. To address these challenges, we build a detector based on the bird’s-eye-view (BEV) detection paradigm, where the explicit feature projection is beneficial to addressing the geometry learning ambiguity when employing multiple scenarios of data to train detectors. Then, we split the classical BEV detection architecture into two stages and propose an uneven BEV grid design to handle the convergence instability caused by the aforementioned challenges. Moreover, we develop a sparse BEV feature projection strategy to reduce computational cost and a unified domain alignment method to handle heterogeneous domains. Combining these techniques, a unified detector UniMODE is derived, which surpasses the previous state-of-the-art on the challenging Omni3D dataset (a large-scale dataset including both indoor and outdoor scenes) by 4.9% $\rm AP_{3D}$, revealing the first successful generalization of a BEV detector to unified 3D object detection.
1 Introduction
Monocular 3D object detection aims to accurately determine the precise 3D bounding boxes of targets using only single images captured by cameras [13, 16]. Compared to 3D object detection based on other modalities such as LiDAR point cloud, the monocular-based solution offers advantages in terms of cost-effectiveness and comprehensive semantic features [19, 17]. Moreover, owing to the wide-ranging applications like autonomous driving [8], monocular 3D object detection has drawn much attention recently.
Thanks to the efforts paid by the research community, numerous detectors have been developed. Some are designed for outdoor scenarios [9, 38] such as urban driving, and the others focus on indoor detection [28]. Despite their common goal of monocular 3D object detection, these detectors exhibit significant differences in their network architectures [5]. This divergence hinders researchers from combining data of various scenarios to train a unified model that performs well in diverse scenes, which is demanded by many important applications like robot navigation [30].
@@ 1 Introduction
Figure 1: Illustration of some challenges (e.g., diverse geometry properties, heterogeneous domain distributions) in unified detection. (1) Comparing sub-figures (a) and (b), indoor objects are small and close, while outdoor objects are far and sparse. Besides, the camera parameters are highly varying. (2) Comparing sub-figures (a), (b), and (c), which correspond to a real-world indoor image, real-world outdoor image, and synthetic indoor image, the image styles are different. (3) Although the category “Picture” is labeled in sub-figure (c), it is not labeled in sub-figure (d), which suggests label conflict among different sub-datasets. Unlabeled objects are highlighted by red ellipses.
@@
The most critical challenge in unified 3D object detection lies in addressing the distinct characteristics of different scenarios. For example, indoor objects are smaller and closer in proximity, while outdoor detection needs to cover a vast perception range. Recently, Cube RCNN [5] has served as a predecessor in studying this problem. It directly produces 3D box predictions in the camera view and adopts a depth decoupling strategy to tackle the domain gap among scenes. However, we observe that it suffers serious convergence difficulty and is prone to collapsing during training.
To overcome the unstable convergence of Cube RCNN, we employ the recent popular bird’s-eye-view (BEV) detection paradigm to develop a unified 3D object detector. This is because the feature projection in the BEV paradigm aligns the image space with the 3D real-world space explicitly [15], which alleviates the learning ambiguity in monocular 3D object detection. Nevertheless, after extensive exploration, we find that naively adopting existing BEV detection architectures [15, 18] does not yield promising performance, which is mainly blamed on the following obstacles.
@@
First of all, as shown in Fig. 1 (a) and (b), the geometry properties (e.g., perception ranges, target positions) between indoor and outdoor scenes are diverse. Specifically, indoor objects are typically a few meters away from the camera, while outdoor targets can be more than 100m away. Since a unified BEV detector needs to recognize objects in all scenarios, the BEV feature has to cover the maximum possible perception range. Meanwhile, as indoor objects are often small, the BEV grid resolution for indoor detection should be precise. All these characteristics lead to unstable convergence and significant computational burden. To address these challenges, we develop a two-stage detection architecture. In this architecture, the first stage produces initial target position estimation, and the second stage locates targets using this estimation as priori, which helps stabilize the convergence process. Moreover, we introduce an uneven BEV grid split strategy that expands the BEV space range while maintaining a manageable BEV grid size. Furthermore, a sparse BEV feature projection strategy is developed to reduce the projection computational cost by 82.6%.
Another obstacle arises from the heterogeneous domain distributions (e.g., image styles, label definitions) across various scenarios. For example, as depicted in Fig. 1 (a), (b), and (c), the data can be collected in real scenes or synthesized virtually. Besides, comparing Fig. 1 (c) and (d), a class of objects may be annotated in a scene but not labeled in another scene, leading to confusion during network convergence. To handle these conflicts, we propose a unified domain alignment technique consisting of two parts, the domain adaptive layer normalization to align features, and the class alignment loss for alleviating label definition conflict.
Combining all these innovative techniques, a Unified Monocular Object DEtector named UniMODE is developed, and it achieves state-of-the-art (SOTA) performance on the Omni3D benchmark. In the unified detection setting, UniMODE surpasses the SOTA detector, Cube RCNN, by an impressive 4.9% in terms of $\rm AP_{3D}$ (average precision based on 3D intersection over union). Furthermore, when evaluated in indoor and outdoor detection settings individually, UniMODE outperforms Cube RCNN by 11.9% and 9.1%, respectively. This work represents a pioneering effort to explore the generalization of BEV detection architectures to unified 3D object detection. It showcases the immense potential of BEV detection across a broad spectrum of scenarios and underscores the versatility of this technology.
2 Related Work
@@
Monocular 3D object detection. Due to its advantages of being economical and flexible, monocular 3D object detection grabs much research attention [22]. Existing detectors can be broadly categorized into two groups, camera-view detectors and BEV detectors. Among them, camera-view detectors generate results in the 2D image plane before converting them into the 3D real space [10, 25]. This group is generally easier to implement. However, the conversion from the 2D camera plane to 3D physical space can introduce additional errors [32], which negatively impact downstream planning tasks typically performed in 3D [7].
@@
BEV detectors, on the other hand, transform image features from the 2D camera plane to the 3D physical space before generating results in 3D [12]. This approach benefits downstream tasks, as planning is also performed in the 3D space [18]. However, the challenge with BEV detectors is that the feature transformation process relies on accurate depth estimation, which can be challenging to achieve with only camera images [23]. As a result, convergence becomes unstable when dealing with diverse data scenarios [5].
@@
Unified object detection. In order to improve the generalization ability of detectors, some works have explored the integration of multiple data sources during model training [34, 14]. For example, in the field of 2D object detection, SMD [40] improves the performance of detectors through learning a unified label space. In the 3D object detection domain, PPT [36] investigates the utilization of extensive 3D point cloud data from diverse datasets for pre-training detectors. In addition, Uni3DETR [35] reveals how to devise a unified point-based 3D object detector that behaves well in different domains. For the camera-based detection track, Cube RCNN [5] serves as the sole predecessor in the study of unified monocular 3D object detection. However, Cube RCNN is plagued by the unstable convergence issue, necessitating further in-depth analysis within this track.
@@
3 Method
3.1 Overall Framework
The overall framework of UniMODE is illustrated in Fig. 2. As shown, a monocular image $I\in\mathbb{R}^{3\times H\times W}$ sampled from multiple scenarios (e.g., indoor and outdoor, real and synthetic, daytime and nighttime) are input to the feature extraction module (including a backbone and a neck) to produce representative feature $F\in\mathbb{R}^{C\times\frac{H}{16}\times\frac{W}{16}}$. Then, $F$ is processed by 4 fully convolutional heads, namely “domain head”, “proposal head”, “feature head”, and “depth head”, respectively. Among them, the role of the domain head is to predict which pre-defined data domain an input image is most relevant to, and the classification confidence produced by the domain head is subsequently utilized in domain alignment. The proposal head aims to estimate the rough target distribution before the 6 Transformer decoders, and the estimated distribution serves as prior information for the second-stage detection. This design alleviates the distribution mismatch between diverse training domains (refer to Section 3.2). The proposal head output is encoded as $M$ proposal queries. In addition, $N$ queries are randomly initialized and concatenated with the proposal queries for the second-stage detection, leading to $M+N$ queries in the second stage.
The feature head and depth head are responsible for projecting the image feature into the BEV plane and obtaining the BEV feature. During this projection, we develop a technique to remove unnecessary projection points, which reduces the computing burden by about 82.6% (refer to Section 3.4). Besides, we propose the uneven BEV feature (refer to Section 3.3), which means the BEV grids closer to the camera enjoy more precise resolution, and the grids farther to the camera cover broader perception areas. This design well balances the grid size contradiction between indoor detection and outdoor detection without extra memory burden.
Obtaining the projected BEV feature, a BEV encoder is employed to further refine the feature, and 6 decoders are adopted to generate the second-stage detection results. As mentioned before, $M+N$ queries are used during this process. After the 6 decoders, the queries are decoded as detection results by querying the FFN. In the decoder part, the unified domain alignment strategy is devised to align the data of various scenarios via both the feature and loss perspectives. Refer to Section 3.5 for more details.
3.2 Two-Stage Detection Architecture
@@
The geometry property difference is identified as an essential reason causing the unstable convergence of BEV detectors [15]. For example, the difference in target position distribution makes it challenging for Transformer-based detectors to learn how to update the query reference points gradually toward the concerned objects. In fact, through visualization, we find the reference point updating in the 6 Transformer decoders is disordered. As a result, if we adopt the classical deformable DETR architecture [41] to build a 3D object detector, the training is easy to collapse due to the inaccurate positions of learned reference points, resulting in sudden gradient vanishing or exploding.
@@
Besides, since the positions of query reference points are not randomly initialized, the iterative bounding box refinement strategy proposed in deformable DETR [41] is abandoned as it may lead to a deterioration of reference points’ quality. In fact, we observe that this iterative bounding box refinement strategy could result in convergence collapse.
3.3 Uneven BEV Grid

### v5

# Towards Unified 3D Object Detection via Algorithm and Data Unification

@@ Abstract
Realizing unified 3D object detection, including both indoor and outdoor scenes, holds great importance in applications like robot navigation. However, involving various scenarios of data to train models poses challenges due to their significantly distinct characteristics, e.g., diverse geometry properties and heterogeneous domain distributions. In this work, we propose to address the challenges from two perspectives, the algorithm perspective and data perspective. In terms of the algorithm perspective, we first build a monocular 3D object detector based on the bird’s-eye-view (BEV) detection paradigm, where the explicit feature projection is beneficial to addressing the geometry learning ambiguity. In this detector, we split the classical BEV detection architecture into two stages and propose an uneven BEV grid design to handle the convergence instability caused by geometry difference between scenarios. Besides, we develop a sparse BEV feature projection strategy to reduce the computational cost and a unified domain alignment method to handle heterogeneous domains. From the data perspective, we propose to incorporate depth information to improve training robustness. Specifically, we build the first unified multi-modal 3D object detection benchmark MM-Omni3D and extend the aforementioned monocular detector to its multi-modal version, which is the first unified multi-modal 3D object detector. We name the designed monocular and multi-modal detectors as UniMODE and MM-UniMODE, respectively. The experimental results reveal several insightful findings highlighting the benefits of multi-modal data and confirm the effectiveness of all the proposed strategies.
Index Terms:
3D Object Detection, Unified Detection, Monocular Detection, Multi-modal Detection Benchmark.
I Introduction
As a fundamental perception task, 3D object detection is to help agents understand the real world. Over the years, it has gained substantial interest from both the academia and industrial communities, leading to the development of numerous 3D object detectors [1, 2, 3, 4, 5]. However, most of these detectors are trained and evaluated within a single scenario, essentially confining them to a limited number of domains. Among the detectors, many are designed for outdoor scenarios such as urban driving [6, 7], and the others focus on indoor detection [8]. In contrast, numerous applications such as robotic navigation, require detectors to perform well in diverse environments, including both indoor and outdoor scenarios [9]. This highlights the need for exploring more unified detectors.
@@ I Introduction
Figure 1: Sub-figures (a)$\sim$(d): Challenges of unified 3D object detection. (1) Comparing sub-figures (a) and (b), indoor objects are small and close, while outdoor objects are far and sparse. Besides, the camera parameters are highly varying. (2) Comparing sub-figures (a), (b), and (c), which correspond to a real-world indoor image, a real-world outdoor image, and a synthetic indoor image, the image styles are different. (3) Although the category “Picture” is labeled in sub-figure (c), it is not labeled in sub-figure (d), which suggests label conflict among different sub-datasets. Unlabeled objects are highlighted by red ellipses. Sub-figures (e)$\sim$(j): Illustrations of the MM-Omni3D benchmark, which showcase both the 3D box annotations and point clouds. The sub-figures clearly demonstrate the significant point cloud differences between different scenarios due to depth sensor discrepancies.
@@
The most critical challenge in unified 3D object detection lies in addressing the distinct characteristics of different scenarios. For example, indoor objects are smaller and closer in proximity, while outdoor detection needs to cover a vast perception range. Recently, Cube RCNN [10] has served as a predecessor in studying this problem. It directly produces 3D box predictions in the camera view and adopts a depth decoupling strategy to tackle the domain gap among scenes. However, we observe that it suffers serious convergence difficulty during training. To overcome the unstable convergence of Cube RCNN, we employ the recent popular bird’s-eye-view (BEV) detection paradigm to develop a unified 3D object detector. This is because the feature projection in the BEV paradigm aligns the image space with the 3D real-world space explicitly [3], which alleviates the learning ambiguity in monocular 3D object detection. Nevertheless, after extensive exploration, we find that naively adopting existing BEV detection architectures [3, 4] does not yield promising performance, which is mainly blamed on the following obstacles.
@@
First of all, as shown in Fig. 1 (a) and (b), the geometry properties (e.g., perception ranges, target positions) between indoor and outdoor scenes are diverse. Specifically, indoor objects are typically a few meters away from the camera, while outdoor targets can be more than 100m away. Since a unified BEV detector needs to recognize objects in all scenarios, the BEV feature has to cover the maximum possible perception range. Meanwhile, as indoor objects are often small, the BEV grid resolution for indoor detection should be precise. All these characteristics lead to unstable convergence and significant computational burden. Another obstacle arises from the heterogeneous domain distributions (e.g., image styles, label definitions) across various scenarios. For example, as depicted in Fig. 1 (a), (b), and (c), the data can be collected in real scenes or synthesized virtually. Besides, comparing Fig. 1 (c) and (d), a class of objects may be annotated in a scene but not labeled in another scene, leading to optimization confusion during the network convergence process.
In this work, we address the challenges of unified 3D object detection from two perspectives, the algorithm perspective and data perspective. From the algorithm perspective, we develop a two-stage BEV detection architecture to tackle the different data geometry properties. In this architecture, the first stage produces initial target position estimation, and the second stage locates targets using this estimation as priori, which helps stabilize the convergence process. Moreover, we introduce an uneven BEV grid split strategy that expands the BEV space range while maintaining a manageable BEV grid size. A sparse BEV feature projection strategy is developed to reduce the projection computational cost by 82.6%. In addition, to handle the heterogeneous domain distributions, we propose a unified domain alignment technique consisting of two parts, the domain adaptive layer normalization (DALN) to align features, and the class alignment loss for alleviating label definition conflict. Combining all these innovative techniques, a well-behaved unified monocular 3D object detector is derived.
In terms of the data perspective, we observe that the unstable training of unified 3D object detectors is partly caused by the ill-posed nature of monocular depth estimation, and this nature can be effectively handled by incorporating depth sensors like lidars, which convert the monocular task into a multi-modal 3D object detection problem. Previous literature has advocated for camera-only detection due to the cost-effectiveness of cameras compared to depth sensors [6]. Nevertheless, the prices of depth sensors are continuously decreasing and have now reached a level applicable for daily use. For instance, all Apple’s iPads and iPhones have been equipped with lidars [11]. Moreover, sparse point clouds are acceptable as deep learning models are capable of inferring depth based on nearby features [12]. Therefore, inexpensive depth sensors present a viable option for multi-modal 3D perception, and we believe exploring unified multi-modal 3D object detection is valuable.
Nevertheless, despite the critical role of data in deep learning, there is no multi-modal 3D object detection benchmark that covers a diverse range of scenarios. To bridge this gap, we release the first multi-modal 3D object detection benchmark MM-Omni3D including multiple indoor and outdoor scenes. Several examples from MM-Omni3D are visualized in Fig. 1 (e)$\sim$(j), where significant variations in image styles and depth point densities can be observed. Each data pair in MM-Omni3D comprises an image and a point cloud, with a total of 159.5K pairs and 629.4K valid 3D box annotations. Afterwards, we extend the aforementioned monocular 3D object detector into its multi-modal version by adding a novel multi-modal feature fusion module named Mutual Information Collaboration (MIC), and it allows the image and point features to complement each other. This multi-modal detector serves as a simple baseline in MM-Omni3D.
In this work, we term the aforementioned monocular and multi-modal unified monocular object detectors as UniMODE and MM-UniMODE, respectively. Extensive experiments are conducted to validate the effectiveness of these two detectors and the developed composite benchmark MM-Omni3D. The results suggest that: (1) UniMODE achieves state-of-the-art (SOTA) performance on Omni3D, which is a large-scale popular monocular 3D object detection dataset. (2) The unstable training dynamics of unified detection can be resolved efficiently by incorporating depth information. (3) Co-training with data from more domains generally improves the detection accuracy and enhances the out-of-domain generalization capability. (4) In addition to serving as a multi-modal 3D object detection benchmark, MM-Omni3D also shows promise as a pre-training dataset.
This paper is an extended version of our conference paper [1], where we make the following new contributions:
$\bullet$  We build the first multi-modal 3D object detection benchmark that covers diverse scenarios, including both indoor and outdoor scenes. This benchmark enables us to study the unified 3D object detection problem not only in the monocular detection setting but also with the multi-modal detection setting.
$\bullet$  We design a novel multi-modal feature fusion module MIC and extend the monocular 3D object detector UniMODE to its multi-modal version, namely MM-UniMODE. This work serves as a pioneering effort to explore the unified multi-modal 3D object detection problem.
$\bullet$  We draw several insightful findings that highlight the benefits of using multi-modal data through performing more extensive experiments. The experimental results indicate that employing multi-modal data for unified 3D object detection can efficiently address the unstable training problem. Besides, we reveal that MM-Omni3D can serve as an effective co-training or pre-training dataset.
II Related Work
@@
Monocular 3D object detection. Due to the advantages of being economical and flexible, monocular 3D object detection grabs much research attention. Existing monocular 3D object detectors can be broadly categorized into two groups, camera-view detectors and BEV detectors. Among them, camera-view detectors first predict attributes that can be represented in the 2D image plane, such as projected 3D centers and depth. Then, the predicted attributes are converted into the 3D physical space based on camera intrinsic and extrinsic parameters [13, 14], and the converted attributes can be used to generate 3D detection boxes. The network designs of camera-view detectors primarily follow the frameworks of popular 2D object detectors. For example, M3D-RPN [15] represents targets as 2D anchor boxes like Faster RCNN [16]. SMOKE [17] utilizes the feature at the positions of the projected 3D centers to describe concerned objects, the idea of which is similar to CenterNet [18]. A significant advantage of this camera-view detection paradigm is that the various attributes composing 3D detection boxes are disentangled, thereby easing the difficulty of analyzing results. However, the conversion from the 2D camera plane to 3D physical space can introduce additional errors [19], which negatively impact downstream planning tasks typically performed in 3D [20].
@@
BEV detectors, on the other hand, transform image feature from the 2D camera plane to the 3D physical space before generating results in 3D [21]. This approach benefits downstream tasks, as planning is also performed in the 3D space [22]. The 2D-to-3D feature transformation strategies adopted by existing BEV detectors can mainly be divided into three classes, i.e., projection based, sampling based, and implicit transformation based. In terms of the projection based strategies, they directly project the camera-view feature to the BEV grids pre-defined in the 3D space [2]. Differently, the sampling based strategies first project the 3D centers of BEV grids back to the camera view to sample feature, and then aggregate the sampled feature to derive BEV feature [3]. For the implicit transformation based strategies, they fuse 2D image feature as 3D feature implicitly through global Transformer attention instead of explicit transformation [4]. Although the three strategies employ different implementations, their motivations are similar, i.e., directly generating detection boxes based on 3D feature. Nevertheless, these strategies also show limitations. All of them require accurate depth estimation (explicit or implicit), which can be challenging to achieve with only camera images [23]. As a result, convergence could become unstable when dealing with diverse data scenarios [10].
Multi-modal 3D object detection. Although some previous detectors solely rely on cameras for cost-effectiveness, their performance is often compromised due to the ill-posed nature of monocular depth estimation [17, 24]. To handle this problem, there are works that exclusively use depth sensors, which yield precise 3D bounding boxes but may struggle with accurate category prediction due to the lack of semantic information in point clouds [25, 26]. As such, multi-modal 3D object detection, which combines cameras and depth sensors, presents a promising approach. The key research problem in multi-modal 3D object detection compared with monocular 3D object detection is how to fuse the multi-modal feature. To study this problem, early methods like EPNet [27] decorate raw depth points with sampled image feature. However, when the input point cloud is sparse, a part of semantic information contained in the image feature could be discarded. There are also methods that utilize input depth points to assist image-based depth estimation and then produce more accurate 3D detection boxes [28]. This paradigm of methods usually depends heavily on precise depth estimation, which is not always attainable because of the sparse input point cloud. Recently, inspired by the BEV detection architectures, several methods first transform the image feature and depth feature into the same space and then fuse them [29]. Compared with previous methods, this paradigm can fully exploit both image and depth information, and the dependence on accurate depth estimation is also eased. Notably, the space used for fusing feature could be implicit. For instance, CMT [30] encodes both image and point feature as tokens and fuses them through iterative Transformer attention based feature interaction. Notably, although there have been plentiful works about multi-modal 3D object detection, current detectors are limited to training and evaluation within a single data scenario, leading to these detectors specializing in a limited number of domains [19]. While the modules adopted by these detectors improve the performance in some domains, they could negatively impact detection precision in others. To bridge this gap, this work expects to develop a multi-modal 3D object detector that is trained and evaluated across multiple data scenarios.
@@
Unified object detection. In order to improve the generalization ability of detectors, some works have explored the integration of multiple data sources during model training [31, 32]. For example, in the field of 2D object detection, SMD [33] observes that a significant challenge existing in unifying multiple domains of data to trained detectors is the label conflict between domains. To address the challenge, this work proposes to use the same detection network but different training protocols for each dataset. In this way, through scaling up the training data volume, a unified label space is learned and the detection accuracy is boosted. Similarly, in the 3D object detection domain, PPT [34] investigates the utilization of extensive 3D point cloud data from diverse datasets for pre-training detectors. Unlike SMD, PPT aligns the knowledge from various domains through training a small set of domain-specific parameters. In addition, Uni3DETR [35] reveals how to devise a unified point-based 3D object detection architecture that can be applied to different datasets. However, an individual detector needs to be trained for each dataset, which means the network parameter is not shared. For the camera-based detection track, Cube RCNN [10] serves as the sole predecessor in the study of unified monocular 3D object detection. However, Cube RCNN is plagued by the unstable convergence issue, necessitating further in-depth analysis within this track.
III Unified Monocular Detection
In this section, we introduce our proposed unified monocular 3D object detector UniMODE. Specifically, in Section III-A, we describe its overall framework. In Section III-B, Section III-C, and Section III-D, we elaborate on the developed two-stage detection architecture, uneven BEV grid, and sparse BEV feature projection strategy, respectively. These designs are to bridge the diverse geometry properties of various training domains. Then, in Section III-E, we explain the strategies utilized to address the heterogeneous domain distributions.
@@ III Unified Monocular Detection
III-A UniMODE Overall Framework
The overall framework of UniMODE is illustrated in Fig. 2. As shown, a monocular image $I\in\mathbb{R}^{3\times H\times W}$ sampled from multiple scenarios (e.g., indoor and outdoor, real and synthetic, daytime and nighttime) are input to the feature extraction module (including a backbone and a neck) to produce representative feature $F\in\mathbb{R}^{C\times\frac{H}{16}\times\frac{W}{16}}$. Then, $F$ is processed by 4 fully convolutional heads, namely “domain head”, “proposal head”, “feature head”, and “depth head”, respectively. Among them, the role of the domain head is to predict which pre-defined data domain an input image is most relevant to, and the classification confidence produced by the domain head is subsequently utilized in domain alignment. The proposal head aims to estimate the rough target distribution before the 6 Transformer decoders, and the estimated distribution serves as prior information for the second-stage detection. This design alleviates the distribution mismatch between diverse training domains (refer to Section III-B). The proposal head output is encoded as $M$ proposal queries. In addition, $N$ queries are randomly initialized and concatenated with the proposal queries for the second-stage detection, leading to $M+N$ queries in the second stage.
The feature head and depth head are responsible for projecting the image feature into the BEV plane and obtaining the BEV feature. During this projection, we develop a technique to remove unnecessary projection points, which reduces the computing burden by about 82.6% (refer to Section III-D). Besides, we propose the uneven BEV feature (refer to Section III-C), which means the BEV grids closer to the camera enjoy more precise resolution, and the grids farther to the camera cover broader perception areas. This design well balances the grid size contradiction between indoor detection and outdoor detection without extra memory burden.
Obtaining the projected BEV feature, a BEV encoder is employed to further refine the feature, and 6 decoders are adopted to generate the second-stage detection results. As mentioned before, $M+N$ queries are used during this process. After the 6 decoders, the queries are decoded as detection results by querying the FFN. In the decoder part, the unified domain alignment strategy is devised to align the data of various scenarios via both the feature and loss perspectives. Refer to Section III-E for more details.
III-B Two-Stage Detection Architecture
@@
The geometry property difference is identified as an essential reason causing the unstable convergence of BEV detectors [3]. For example, the difference in target position distribution makes it challenging for Transformer-based detectors to learn how to update the query reference points gradually towards the concerned objects. In fact, through visualization, we find the reference point updating in the 6 Transformer decoders is disordered. As a result, if we adopt the classical deformable DETR architecture [36] to build a 3D object detector, the training is easy to collapse due to the inaccurate positions of learned reference points, resulting in sudden gradient vanishing or exploding.
@@
Besides, since the positions of query reference points are not randomly initialized, the iterative bounding box refinement strategy proposed in deformable DETR [36] is abandoned as it may lead to a deterioration of reference point quality. In fact, we observe that this iterative bounding box refinement strategy could result in convergence collapse.
III-C Uneven BEV Grid